The combined ages of a father and his twin daughters are 54 years.The father was 24 years when the twins were born.How old is each of the three people?
Is it the father was 24 years,and each twin was 15 years because
54-24=30
30/2=15
or is it 54-24=30,30/2=15,24+15=39 years old for the father and 15 years old for the twins each?

Answer by scott8148(6628)

f + 2t = 54

f = t + 24

substituting ___ (t + 24) + 2t = 54 ___ t = 10